22. What is the pH of 0.12 M HCIO(aq) (Ka=3.0 x 10-)?
Answer #1

HA (aq) + H₂O (1) Ž H₂ot (al + A Cow 0 x 0 x (Initial) (Equilibrium) (a-xl concentration of at Initial Molarity. let cx be ta-0.121 39168 (0-12-X) X*- 391020-12 - 3x10-87- 22 + (3x16-37 ) - ( 3x10 - 0-12) = 0 2 - 0-00005 ११85 CH, ) = 0-000 Sq4 85 PH